# references.bib

### v2

@@ Abstract
Early detection of fertilizer-induced stress in tomato plants is crucial for timely crop management interventions and yield optimization. Conventional optical methods detect fertilizer stress in young leaves with difficulty. This study proposes a novel, noninvasive technique for quantifying the density of trichomes-elongated hair-like structures found on plant surfaces-on young leaves using a smartphone. This method exhibits superior detection latency, enabling earlier and more accurate identification of fertilizer stress in tomato plants. Our approach combines augmented reality technology and image processing algorithms to analyze smartphone images of a specialized measurement paper. This measurement paper is applied to a tomato leaf to transfer trichomes onto its adhesive surface. The captured images are then processed through a pipeline involving region of interest extraction, perspective transformation, and illumination correction. Trichome detection and spatial distribution analysis of these preprocessed images yield a robust density metric. We validated our method through experiments on hydroponically grown tomatoes under varying fertilizer concentrations. Using leave-one-out cross-validation (LOOCV), our model achieves a mean area under the precision-recall curve of 0.824 and a receiver operating characteristic curve of 0.641 for predicting additional fertilization needs. Based on LOOCV, quantitative analysis revealed a strong relationship between trichome density and explanatory variables, including nitrate ion concentration, explaining 62.48% of the variation ($R^{2}=0.625$). The predicted and actual trichome densities were strongly correlated ($r=0.794$). This straightforward and cost-effective method overcomes the limitations of traditional techniques, demonstrating the potential of using smartphones for practical plant nutrition diagnosis.
@@ 1 Introduction
Tomato (Solanum lycopersicum), a vegetable that is widely cultivated worldwide, is influenced by various environmental factors that affect its production and quality. The nutrient status of plants, one of the most crucial factors influencing tomato yield, is an important consideration in tomato production. Therefore, accurately assessing plant nutrient status is essential for maintaining high tomato yields. Techniques have been developed to analyze the constituents of fertilizers in soil, media, and nutrient solutions, aiming to estimate the amount of nutrients absorbed by plants [sonneveld1990estimating, chen2015crop] . Additionally, researchers have developed invasive methods to analyze plant nutrients [hochmuth1994efficiency, andersen1999relationships] . Moreover, noninvasive techniques such as visual inspection and sensor-based monitoring are available for assessing plant traits.
@@
Morphological changes can serve as valuable indicators for nutritional analysis, particularly in the context of environmental stressors such as drought and nutrient deficiencies. With respect to drought stress, Klepper et al. conducted an observational study and reported that cotton stem diameter, leaf water potential, and relative leaf water content are closely related to the net radiation above the plant canopy [klepper1971stem]. Wakamori et al. optically analyzed the wilting of tomato leaves and changes in stem diameter in an observational study and demonstrated a correlation between these parameters[wakamori2019optical]. In terms of nutrient stress, Padilla et al. provided a comprehensive review of techniques for monitoring the nitrogen status of vegetable crops via optical methods [padilla2018proximal]. Locascio et al. and Coltman demonstrated a causal relationship between the fertilization rate and tomato yield in interventional studies[locascio1997nitrogen, coltman1988yields]. Tei et al. reported a causal relationship between the fertilization rate and tomato leaf area in an interventional study[tei2002critical]. These findings emphasize that environmental stressors alter the vegetative and reproductive growth patterns observed in plants.
@@
The ability to detect and examine these modifications allows for a more sophisticated understanding of the nutrient levels of plants and their responses to varying environmental conditions. Leaf area and chlorophyll content, which are related to leaf color and spectrum, are commonly used as indicators of the adaptive response of plants to fertilizer stress. dela Torre et al. reviewed various optical indices, leaf area, biomass, and other phenotypes for estimating rice yield via remote sensing[delatorre2021remote]. Cartelat et al. demonstrated that optically assessed contents of leaf polyphenols and chlorophyll in wheat can be used as indicators of nitrogen deficiency[cartelat2005optically]. Padilla et al. reported that chlorophyll meters and canopy reflectance sensors serve as sensitive indicators of N status across a wide range of vegetable crops[padilla2020monitoring]. These methods have led to the development of a diagnostic framework based on the relationships among fertilizer application, phenotypic expression, and yield outcomes. Color and spectral data analysis of plant canopies enables predictions of plant health and productivity. These predictions utilize wide-area imagery from satellites, aircraft, and drones combined with data collected from in-field sampling stations. The aforementioned optical observation techniques, which utilize reflected light from the plant body, effectively address the need for noninvasive and digital monitoring of phenotypes [padilla2018proximal].
@@
Furthermore, techniques for detecting fertilizer stress in tomatoes have been extensively researched. Tables 1 and 2 present a comprehensive comparison of various noninvasive methods for early detection of fertilizer stress in tomato plants. We focused on studies that employed interventional experiments with varying fertilizer levels and used noninvasive measurement techniques on tomatoes. Priority was given to methods that utilized digital measurement tools and enabled measurements as close to the growth point as possible for early stress detection. As shown in Table 1, although these methods may appear equally effective at first glance, the complexity of data processing (visualization, modeling, and interpretation) after measurement must be considered, which varies considerably depending on the selected phenotypic indicator and sensor type.
@@
Moreover, in terms of detection latency, which refers to the delay between the onset of fertilizer stress and its detection, evaluating phenotypic indicators based on the metabolic mechanisms of tomato plants is crucial. Tomatoes exhibit apical growth, where cell differentiation occurs at the top of the plant, and differentiated organs move downward as the plant grows. Considering this mechanism, measuring phenotypic indicators near the shoot apical meristem could lead to earlier and more accurate detection of fertilizer stress. However, as shown in Table 2, existing studies primarily utilize "fully expanded leaves" or even the entire plant as a phenotypic indicator for chlorophyll meters, visible-near infrared spectroscopy, fluorescence spectroscopy, stem diameter measurement, and leaf area measurement. A "fully expanded leaf" is generally considered the 3rd to 5th compound leaf from the apex, excluding the apical meristem and leaf primordia, and has reached its final size and shape. However, a universally accepted definition of a fully expanded leaf does not exist, and its specific characteristics are as follows: The precise distance from the apical meristem may vary depending on factors such as plant species, cultivar, growth stage, and environmental conditions. Therefore, the context of each study must be considered when interpreting the term "fully expanded leaf." Consequently, studies in the literature identifying the phenomenon of fertilizer stress are lacking. This scarcity is especially true in the assessment of young tomato leaves, with the definition of "young leaves" itself being subject to variation.
@@
* This table includes studies that met the following criteria: noninvasive measurement, tested on tomatoes, interventional experiments with varying fertilizer levels. Priority was given to methods with digital measurements and those enabling measurements near the growth point for early stress detection.
@@
a Measurement location not specified in the original study.
@@
d mPR and mROC represent the mean area under the precision-recall curve and the mean area under the receiver operating characteristic curve, respectively (see equations 10 and 11). These metrics evaluate the performance of the model in determining the need for additional fertilization.
@@ 2 Related Work
Gianquinto et al. conducted a study using a multispectral radiometer positioned directly above the tomato plant canopy to investigate the correlation between light reflectance, chlorophyll concentration, and total Kjeldahl nitrogen content in an intervention study of fertilizer application. By employing this noncontact arrangement, measurements were conducted without physically interacting with the leaves. They discovered a linear relationship between optical reflectance and chlorophyll concentration; however, no such linear relationship was identified between the optical reflectance and total Kjeldahl nitrogen content [gianquinto2011a]. Similarly, Padilla et al. directly observed the most recently fully expanded leaves via an optical leaf-clip chlorophyll meter (Konica Minolta, SPAD-502). A linear relationship was observed between optical observations and the nitrogen nutrition index under conditions of a flattened leaf surface structure [padilla2015threshold]. In both investigations, the absence of any barriers along the path of light between the observing device and the plant leaves is assumed, highlighting a gap between the specifications of diagnostic equipment for farmers and the current state of research and development.
However, in contrast to other crops, tomatoes have numerous trichomes, which are small projections or hair-like structures, on the surface of their leaves (Figure 1). Trichomes are densely concentrated on young, actively growing leaves and sparsely distributed on older leaves [johnson1975plant]. The presence of surface microstructures, such as trichomes, on plant bodies is known to considerably affect their appearance [kinoshita2005structural, middleton2024selfassembled, thomas2010function]. Given their high density on young tomato leaves, it is plausible that trichomes may influence the effectiveness of optical sensing techniques used to assess the nutrient status of tomato plants. Consequently, the presence of trichomes should be considered a potential practical limitation when optical sensing techniques are used on young tomato leaves to evaluate fertilizer stress. However, current methodologies do not adequately address the impact of trichomes on the optical sensing techniques used for the rapid assessment of tomato plant nutrient status.
@@
Counting trichomes, which are abundant on young tomato leaves, is not easily achievable via visual means. Manually counting trichomes via a microscope is one of the few methods available for quantifying their density. Trichomes are abundant on young tomato leaves and are not easily distinguishable by visual means alone. The lack of a specialized device for assessing trichome density may have hindered progress in this area of research.
The chlorophyll content, which is correlated with leaf color, is a widely used indicator of fertilizer stress in plants. However, it is not the only marker. Trichomes have also been identified as potential indicators of fertilizer stress, as they contribute to plant defense against pests by responding to fertilizer stress and altering plant density [barbour1991interaction, hoffland2000nitrogen]. Various types of trichomes can be observed on the surface of tomato leaves, most of which are short, hair-like trichomes and trichomes with large, spherical cells at their tips (known as type VI trichomes, Figure 2) [wilkens1996resource]. The glandular head of type VI trichomes is connected to the stalk cell at the base by fragile cell walls [bergau2015the].
@@
This structural arrangement allows the glandular part to easily detach from the leaf surface upon physical stimulation. The mucilage within the glandular part is known to inhibit pest activity.
@@
Researchers have observed multiple instances where trichomes respond to fertilizer stress [barbour1991interaction, hoffland2000nitrogen, li2014the]. For example, Barbour et al. revealed an inverse relationship between fertilizer application, leaf nitrogen content, and trichome density [barbour1991interaction]. Similarly, Hoffland et al. reported a positive correlation between the leaf carbon:nitrogen ratio and trichome density [hoffland2000nitrogen]. In contrast, Leite et al. presented data indicating that trichome density increased over time when less fertilizer was used compared with that of conventional farming methods [leite]. Despite these findings, most previous studies have focused on investigating trichome density within a biological context, with limited exploration of techniques to observe trichome density in agricultural production settings without advanced laboratory equipment. This limitation highlights the need for further research to develop practical methods for assessing trichome density in real-world agricultural environments.
To address this gap, the present study focused on tomato leaf trichomes and explored their potential as indicators for assessing the nutrient status of tomato plants. As depicted in Figure 3, our main hypothesis is that leaf trichome density is influenced by fertilizer level. These findings suggest its potential as an additional indicator of plant nutrient status in conjunction with traditional fertilizer stress markers such as the leaf nitrate content. To this end, a tool for assessing trichome density must be developed. The identification of distinct items within a particular setting containing several objects inside the visual field is commonly considered a computer vision problem [zhao2019object, vincent1991watersheds]. Given this context, we hypothesized that the use of cameras could convert the challenge of measuring trichome density to an analogy of an object detection problem.
@@
Figure 3: Hypothesized relationships between fertilizer level, cell differentiation, leaf nitrate content, number of leaf trichomes, and crop yield for tomato plants, serving as the conceptual framework for the study.
@@
Recently, the increasing prevalence of smartphones, with their advanced computational capabilities, high-resolution cameras, and internet access, has spurred the development of proof-of-concept agricultural technologies. These technologies have a wide range of applications, including farm management, crop monitoring, and pest and disease diagnosis, as highlighted in a comprehensive review by Mendes et al. [mendes2020smartphone]. For example, Neumann et al. demonstrated a method for classifying diseases in sugar beet leaves using images captured directly with a smartphone camera [neumann20142014], and Johannes et al. similarly classified wheat diseases via direct leaf images [elhassouny20192019]. Furthermore, Li et al. successfully achieved early detection of late blight in tomato leaves using measurement sheet[li2019noninvasive]. In their study, images of a measurement sheet used to analyze leaf volatiles were captured via a smartphone. Inspired by these advancements, our study aimed to investigate the feasibility of using smartphones as a cost-effective tool for measuring trichome density. These parameters can serve as predictive indicators of tomato plant nutrient status.

### v3

@@ Abstract
Early detection of fertilizer-induced stress in tomato plants is crucial for optimizing crop yield through timely management interventions. While conventional optical methods struggle to detect fertilizer stress in young leaves, these leaves contain valuable diagnostic information through their microscopic hair-like structures, particularly trichomes, which existing approaches have overlooked. This study introduces a smartphone-based noninvasive technique that leverages mobile computing and digital imaging capabilities to quantify trichome density on young leaves with superior detection latency. Our method uniquely combines augmented reality technology with image processing algorithms to analyze trichomes transferred onto specialized measurement paper. A robust automated pipeline processes these images through region extraction, perspective transformation, and illumination correction to precisely quantify trichome density. Validation experiments on hydroponically grown tomatoes under varying fertilizer conditions demonstrated the method’s effectiveness. Leave-one-out cross-validation revealed strong predictive performance with the area under the precision-recall curve (PR-AUC: 0.82) and area under the receiver operating characteristic curve (ROC-AUC: 0.64), while the predicted and observed trichome densities exhibited high correlation ($r=0.79$). This innovative approach transforms smartphones into precise diagnostic tools for plant nutrition assessment, offering a practical, cost-effective solution for precision agriculture.
@@ 1 Introduction
Tomato (Solanum lycopersicum), a vegetable that is widely cultivated worldwide, is influenced by various environmental factors that affect its production and quality. The nutrient status of plants, which is one of the most crucial factors influencing tomato yield, is an important consideration in tomato production. Therefore, accurately assessing plant nutrient status is essential for maintaining high tomato yields.
The assessment of plant nutrient status has evolved from invasive to noninvasive approaches. Initially, techniques were developed to analyze the constituents of fertilizers in soil, media, and nutrient solutions, aiming to estimate the amount of nutrients absorbed by plants [sonneveld1990estimating, chen2015crop]. Additionally, researchers have developed invasive methods to analyze plant nutrients [hochmuth1994efficiency, andersen1999relationships]. More recently, noninvasive techniques, including visual inspection and sensor-based monitoring, have become available for assessing plant traits, offering the advantage of continuous monitoring without damaging the plant.
@@
Among these noninvasive techniques, morphological observation has proven particularly valuable for assessing plant responses to environmental stressors, serving as valuable outcome measures of plant stress responses. Studies have demonstrated strong relationships between morphological changes and various stress conditions, particularly in the context of environmental stressors such as drought and nutrient deficiencies. Under drought stress, Klepper et al. conducted an observational study and reported that cotton stem diameter, leaf water potential, and relative leaf water content are closely related to the net radiation above the plant canopy [klepper1971stem]. Wakamori et al. optically analyzed the wilting of tomato leaves and changes in stem diameter in an observational study and demonstrated a correlation between these parameters [wakamori2019optical]. In terms of fertilizer stress, Padilla et al. provided a comprehensive review of techniques for monitoring the nitrogen status of vegetable crops via optical methods [padilla2018proximal]. Locascio et al. and Coltman demonstrated a causal relationship between the fertilization rate and tomato yield in interventional studies [locascio1997nitrogen, coltman1988yields]. Tei et al. reported a causal relationship between the fertilization rate and tomato leaf area in an interventional study [tei2002critical]. These findings collectively demonstrate how environmental stressors manifest in observable changes in plant morphology.
@@
The ability to detect and examine these modifications allows for a more sophisticated understanding of plant responses to varying environmental conditions. Building on these morphological indicators, optical measurement techniques have been developed to provide quantitative assessments of plant stress responses. These methods focus particularly on leaf characteristics such as area and chlorophyll content, which are commonly used as indicators of the adaptive response of plants to fertilizer stress and can be measured through leaf color and spectral analysis. dela Torre et al. reviewed various optical indices, leaf area, biomass, and other phenotypes for estimating rice yield via remote sensing [delatorre2021remote]. Cartelat et al. demonstrated that optically assessed contents of leaf polyphenols and chlorophyll in wheat can be used as indicators of nitrogen deficiency [cartelat2005optically]. Padilla et al. reported that chlorophyll meters and canopy reflectance sensors serve as sensitive indicators of N status across a wide range of vegetable crops [padilla2020monitoring]. These methods have led to the development of a diagnostic framework based on the relationships among fertilizer application, phenotypic expression, and yield outcomes. Color and spectral data analysis of plant canopies makes it possible to predict plant health and productivity.
@@
Current optical observation techniques utilize various platforms, from handheld devices to satellites and drones, combined with data collected from in-field sampling stations. These techniques, which utilize reflected light from the plant body, enable noninvasive and digital monitoring of plant phenotypes [padilla2018proximal]. The detection of fertilizer stress in tomatoes has been extensively researched via these approaches. Tables 1 and 2 present a comprehensive comparison of various noninvasive methods for the early detection of fertilizer stress in tomato plants. We focused on studies that employed interventional experiments with varying fertilizer levels and used noninvasive measurement techniques on tomatoes. Priority was given to methods that utilized digital measurement tools and enabled measurements as close to the growth point as possible for early stress detection. However, the effectiveness of these methods varies considerably in their practical application. The selected phenotypic indicator and sensor type influence multiple aspects of measurement. These aspects include the complexity of data processing for visualization, modeling, and interpretation; the necessary prerequisites during measurement; and practical constraints such as measurement time and operational complexity.
A critical consideration in tomato nutrient assessment is detection latency, the delay between stress onset and detection. Evaluating phenotypic indicators based on the metabolic mechanisms of tomato plants is crucial, as tomatoes exhibit apical growth, where cell differentiation occurs at the top of the plant and differentiated organs move downward as the plant grows. Considering this mechanism, measuring phenotypic indicators near the shoot apical meristem could theoretically provide earlier and more accurate detection of fertilizer stress. The measurement of young leaves should therefore be recognized as a critical nonfunctional requirement for early stress detection.
@@
However, as shown in Table 2, existing studies primarily utilize "fully expanded leaves" or even the entire plant as a phenotypic indicator for chlorophyll meters, visible-near infrared spectroscopy, fluorescence spectroscopy, stem diameter measurement, and leaf area measurement. A "fully expanded leaf" is typically considered the 3rd to 5th compound leaf from the apex, excluding the apical meristem and leaf primordia, and has reached its final size and shape. Nevertheless, there is no universally accepted definition of a "fully expanded leaf," and its specific characteristics, including the precise distance from the apical meristem, may vary depending on factors such as plant species, cultivar, growth stage, and environmental conditions. Therefore, it is essential to consider the context of each study when interpreting the term "fully expanded leaf".
This approach may delay stress detection, highlighting a significant gap in early stress detection capabilities. There is a lack of studies identifying the phenomenon of fertilizer stress, particularly in the assessment of young tomato leaves, with the definition of "young leaves" itself being subject to variation. Although young leaf measurement is a critical nonfunctional requirement, current methodologies have not adequately addressed this need. To address these limitations and develop more effective methods for early stress detection, a systematic review of existing measurement approaches and their limitations is necessary, particularly with a focus on the challenges specific to tomato plants.
@@
* This table includes studies that met the following criteria: noninvasive measurement, tested on tomatoes, and interventional experiments with varying fertilizer levels. Priority was given to methods with digital measurements and those enabling measurements near the growth point for early stress detection.
@@
a Measurement location was not specified in the original study.
@@
d mPR and mROC represent the mean area under the precision-recall curve and the mean area under the receiver operating characteristic curve, respectively (see equations 12 and 13). These metrics evaluate the performance of the model in determining the need for additional fertilization.
@@ 2 Related Work
To address the challenges identified in early stress detection, particularly in young tomato leaves, this section reviews optical measurements that have emerged as promising approaches for detecting fertilizer stress. In our literature review, we specifically focused on studies that examined plant responses under varying fertilizer conditions, whether through controlled interventions or through observations across fields with distinct fertilization regimes. We distinguished these studies from studies conducted under uniform fertilizer conditions that only identified correlations between chemical analyses and plant phenotypes.
Previous studies have explored various optical methods for assessing plant nutrient status. Of particular interest were studies that combined controlled fertilizer treatments with noninvasive optical assessment methods, as they provide insights into both measurement methodology and its practical limitations. For example, Gianquinto et al. conducted an intervention study in which a multispectral radiometer positioned directly above the tomato plant canopy was used to investigate the correlations among light reflectance, chlorophyll concentration, and total Kjeldahl nitrogen content. By employing this noncontact arrangement, measurements were conducted without physically interacting with the leaves. They discovered a linear relationship between the optical reflectance and chlorophyll concentration; however, no such linear relationship between the optical reflectance and total Kjeldahl nitrogen content was identified [gianquinto2011a]. Similarly, Padilla et al. directly observed the most recently fully expanded leaves via an optical leaf-clip chlorophyll meter (Konica Minolta, SPAD-502). A linear relationship was observed between optical observations and the nitrogen nutrition index under conditions where the leaf surface was mechanically flattened by the measurement clips [padilla2015threshold]. In both investigations, the absence of any barriers along the path of light between the observing device and the plant leaves was assumed, highlighting a gap between the specifications of diagnostic equipment for farmers and the current state of research and development.
While these optical methods have shown promise in various crops, tomato leaves present unique characteristics that create both challenges and opportunities for nutrient status assessment. Unlike many other crops, tomatoes have numerous trichomes, which are small projections or hair-like structures, on the surface of their leaves (Figure 1). These trichomes are densely concentrated on young, actively growing leaves and sparsely distributed on older leaves [johnson1975plant]. The presence of such surface microstructures is known to considerably affect their appearance [kinoshita2005structural, middleton2024selfassembled, thomas2010function]. Given their high density on young tomato leaves, it is plausible that trichomes significantly influence the effectiveness of optical sensing techniques used to assess plant nutrient status. Current methodologies do not adequately address this impact when evaluating fertilizer stress, particularly in young leaves. Therefore, the presence of trichomes should be recognized as an overlooked nonfunctional requirement in measurement device design, and it must be incorporated into the design process.
@@
However, trichomes themselves may offer a novel approach for stress detection. In addition to posing a measurement challenge, trichomes have been identified as potential indicators of fertilizer stress, as they contribute to plant defense against pests by responding to fertilizer stress and altering plant density [barbour1991interaction, hoffland2000nitrogen]. Various types of trichomes can be observed on the surface of tomato leaves, most of which are short, hair-like trichomes and trichomes with large, spherical cells at their tips (known as type VI trichomes; Figure 2) [wilkens1996resource]. The glandular head of type VI trichomes is connected to the stalk cell at the base by fragile cell walls [bergau2015the]. This structural arrangement allows the glandular part to easily detach from the leaf surface upon physical stimulation. The mucilage within the glandular part is known to inhibit pest activity. This detachment mechanism, along with the mucilage that inhibits pest activity, demonstrates how trichomes actively respond to environmental conditions. While chlorophyll content, which is correlated with leaf color, remains a widely used indicator of fertilizer stress in plants, the structural and functional characteristics of trichomes suggest their potential as alternative stress indicators.
@@
Multiple studies have demonstrated relationships between trichomes and nutrient status. Researchers have observed multiple instances where trichomes respond to fertilizer stress [barbour1991interaction, hoffland2000nitrogen, li2014the]. For example, Barbour et al. revealed an inverse relationship between fertilizer application, leaf nitrogen content, and trichome density [barbour1991interaction]. Similarly, Hoffland et al. reported a positive correlation between the leaf C:N ratio and trichome density [hoffland2000nitrogen]. In contrast, Leite et al. reported that trichome density increased over time under reduced fertilizer conditions compared to conventional farming methods [leite]. However, these studies primarily investigated trichome density within a biological context, with limited exploration of practical measurement techniques for agricultural settings [barbour1991interaction, hoffland2000nitrogen, li2014the].
Based on these methodological limitations, we identified two potential approaches for overcoming the challenge of early stress detection in young tomato leaves. These approaches involve modifying existing optical methods through additional data processing and fundamentally redesigning the measurement strategy to utilize trichomes as indicators. To properly account for trichomes in optical measurements, methods for quantifying their density are essential, regardless of which approach is chosen. The latter approach is particularly promising because it could satisfy both the prerequisites for optical measurements and the nonfunctional requirement of early stress detection through young leaf measurements with simpler implementation, as it directly utilizes the surface structures that are inherently abundant in young leaves. As illustrated in Figure 3, our main hypothesis derived from the literature review is that leaf trichome density is influenced by the fertilizer level and could serve as an additional indicator of plant nutrient status, complementing traditional outcome measures such as leaf nitrate content and fruit yield. To address the gap noted above, the present study focused on tomato leaf trichomes and explored their potential as indicators for assessing the nutrient status of tomato plants in real-world agricultural environments.
@@
Figure 3: Hypothesized relationships between the fertilizer level, cell differentiation, leaf nitrate content, the number of leaf trichomes, and crop yield for tomato plants, serving as the conceptual framework for the study.
To implement this approach, however, practical methods for quantifying trichome density must be developed. Currently, manual counting via microscopy remains one of the few available methods, and the lack of specialized measurement tools may have hindered progress in this area. The identification of distinct items within a particular setting containing several objects inside the visual field is commonly considered a computer vision problem [zhao2019object, vincent1991watersheds]. Given this context, we hypothesized that the use of cameras could transform the challenge of measuring trichome density into an analogy of an object detection problem.
@@
Given these challenges in trichome measurement and the need for practical field-based solutions, we explored the potential of widely available technology. As highlighted in a comprehensive review by Mendes et al. [mendes2020smartphone], with their portability, advanced computational capabilities, high-resolution cameras, and internet connectivity, smartphones have enabled various agricultural monitoring applications, including farm management, crop monitoring, and pest and disease diagnosis. Several successful implementations demonstrate the potential of smartphone-based imaging. For example, Neumann et al. developed methods for classifying sugar beet leaf diseases [neumann20142014], Elhassouny et al. obtained similar results with tomato diseases [elhassouny20192019], and Li et al. demonstrated early detection of tomato late blight via smartphone images of measurement sheets [li2019noninvasive]. Building on these successes in detecting fine-scale plant features, our study investigated the feasibility of using smartphones as a cost-effective tool for measuring trichome density as an indicator of tomato plant nutrient status.